OpenStudy (anonymous):

Victoria has two porcelain tea sets that are similar in size. The larger teapot is 12 inches tall and 6 inches wide. If the smaller teapot is 8 inches tall, how wide is it?

OpenStudy (anonymous):

since they are similar their dimensions are proportional

OpenStudy (anonymous):

8/12 = 2/3 so the required width will be 6* 2/3 = 4